1x-2y+3z=37
-2x-3y=2z=16
-2x-1y+1z=-2
solve for x y and z

Answer by Alan3354(21610)

+1x-2y+3z=37
-2x-3y=2z=16 Too many equal signs
-2x-1y+1z=-2
------------